The sum of three consecutive even integers is 210. Find the integers.
Answer by mananth(16949)

let the integers be x,x+2,x+4
..
Their sum = 210
3x+6 = 210
-6
3x=210-6
3x=204
/3
x= 68
...
68,70,72
...
m.ananth@hotmail.ca